# Towards Verifying Exact Conditions of Density Functional Theory Approximations

### v1

@@ Abstract
This paper presents XCVerifier, the first approach for verifying whether a DFA implementation satisfies the DFT exact conditions. XCVerifier was evaluated on five DFAs from the popular Libxc library and seven exact conditions used in recent work by Pederson and Burke. XCVerifier was able to verify or find violations for a majority of the DFA/condition pairs, demonstrating the feasibility of using formal methods to verify DFA implementations. However, it timed out on all conditions of the recent SCAN functional, revealing potential directions for future work.
@@ I Introduction
Density functional theory (DFT) [1, 2, 3] is a widely used approximation in the computation of electronic properties of matter. DFT calculations have been used to predict diverse properties–from atomic binding energies and chemical reactivities to electronic conductivities and magnetic phenomena [4]. Consequently, DFT finds applications in a number of scientific and engineering fields, including solid state physics, computational chemistry, and materials science, and has been implemented in widely-used scientific software such as ABINIT [5], cp2k [6], ERKALE [7], Psi4 [8], Octopus [9], Qbox [10], and Quantum ESPRESSO [11].
Originating in the foundational work of Hohenberg and Kohn in 1964 [12], DFT has provided a practical way to reduce the complexity of conventional electronic structure methods by identifying the electron density function $n(r)$—a real function in three-dimensional space—as the fundamental quantity from which all other properties of a physical system can be derived. The Kohn-Sham (KS) [13] approach is currently the basis of most DFT calculations. The KS DFT approach states that only the exchange-correlation (XC) energy portion of a functional needs to be approximated. However, the exact expression for this XC energy functional $E_{\rm xc}[n]$, which describes the complex electron-electron interactions within systems, is not known and is incredibly difficult to approximate [14].
@@
This proliferation of DFAs is reflected in the fact that the popular Libxc software library [18], which provides numerical implementations of DFAs, currently includes over 500 functionals [19]. The DFAs are also of varying degrees of complexity [20]: (i) local density approximations (LDAs) depend only on the electron density $n$, (ii) generalized-gradient approximations (GGAs) [21] depend on the electron density $n$ and its gradient $\nabla n$, and (iii) meta-GGAs have a further dependence on the Laplacian $\nabla^{2}n$ and the local kinetic energy density $\tau$.
@@
$\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}(\rho,\sigma)=\frac{2.884\rho^{1/3}\left(-28.944\pi^{4/3}\rho^{8/3}-0.174\pi^{2}\sigma\right)}{\pi^{1/3}\cdot\left(77.184\pi^{4/3}\rho^{8/3}+0.257\pi^{2}\sigma\right).}$
@@
However, the correlation part of PBE is significantly more complex with over 300 operations in the Libxc implementation. The SCAN meta-GGA functional [24] is even more complex with over 1000 operations in the exchange and correlation parts combined, including transcendental functions such as $\exp$ and $\log$.
@@
Creating a new DFA is an art that has only been mastered by few researchers as of today. The functional forms used in the definition of a DFA vary considerably. DFA designs fall into two categories: empirical and non-empirical. Empirically-designed DFAs (e.g., LYP [25]) are tailored for molecular chemistry applications and perform well on molecular benchmarks [26, 27]. Non-empirically designed DFAs are constructed to satisfy some exact conditions, which are known analytical properties of the exact functional (Section II). For example, the correlation energy $E_{C}$ non-positivity condition states that the correlation energy cannot be positive, i.e., $E_{\rm c}[n]\leq 0$ [28]. Furthermore, so-called norms are imposed on DFAs by requiring that they reproduce correctly some known physical systems, e.g. a hydrogen or a helium atom for which exact results are available. The SCAN functional is built to satisfy as many as 17 constraints and norms [24].
@@
Recently, Pederson and Burke (PB) [28] checked whether the Libxc implementations of various DFAs satisfy DFT exact conditions. In particular, they checked seven exact conditions by considering their corresponding local condition (Section II). These local conditions were assessed for a given DFA by performing a grid search over the input variables, and checking whether each input-output pair satisfies the local condition. The PB approach is the state of the art in the DFT community, and was the first to perform a large scale study of the role of exact conditions in density functional development. For instance, they found that many empirical DFAs satisfy these exact conditions in certain regions even though they were designed without explicit adherence to these exact conditions. Many of the conditions require gradient or limit calculations, which need to be numerically approximated.
This paper addresses the problem of automatically verifying whether the implementation of a DFA satisfies the exact conditions of the density functional. It is the first to apply formal-methods techniques to density functional theory. The aim is to provide formal guarantees related to the correctness of existing DFAs; viz., to formally verify if the DFA satisfies the exact conditions, and to determine the input regions in which it does not. The prior PB approach might miss violations of the property (because of the use of sampling) or report violations are not true (because many local conditions require gradient or limit calculations that were numerically approximated).
As a step towards solving this problem, we designed and implemented XCVerifier, a tool that automatically converts a Libxc functional (implemented in Maple[29]) to Python, uses dReal [30] and SymPy [31] to compute any required derivatives and limits in the local condition, and then uses the dReal solver to verify that the DFA satisfies the condition in the relevant input region (Section III). We also implemented a domain-splitting technique to improve the performance of XCVerifier (Algorithm 1).
@@
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke) for five popular DFAs: PBE [22], SCAN [24], LYP [25], AM05 [32], and VWN RPA [33], which cover the different types of DFAs (LDA, GGA, and meta-GGA), as well as different design categories (empirical and non-empirical). Some conditions do not apply to certain DFAs, which left us with 29 DFA/condition pairs. As shown in Section IV-B, XCVerifier was successfully able to verify or find a counterexample for 11 pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs (Table I). These results demonstrate the feasibility of using formal-methods techniques to verify DFA implementations. However, XCVerifier timed out for 11 pairs: it was unable to verify the $U_{\text{c}}$ monotonicity for all but the simplest of the DFAs we considered (VWN RPA), and was unable to verify any of the properties for the SCAN DFA. This motivates further research on improving formal-methods techniques in the context of DFT, an important scientific computation discipline.
We also compared the results from the PB approach with XCVerifier (Section IV-C). We found that the results of the PB approach were inconsistent with those of XCVerifier for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions (Table II). For instance, the PB approach found no violations to the $T_{\text{c}}$ upper bound condition for the PBE functional (Figure 2(b)), while XCVerifier was able to find a number of violations (Figure 2(e)). The $T_{\text{c}}$ upper bound condition entails computing derivatives and limits for which the PB approach uses numerical approximations. In contrast, XCVerifier computes these symbolically. This demonstrates one advantage of our approach to verifying DFT exact conditions.
@@
Our preliminary results demonstrate the feasibility of using formal methods to improve the correctness of DFT implementations, and reveal potential avenues for future work (Section VI).
@@ II Exact Conditions in DFT
This section lists the exact conditions of the density functional $E_{\rm xc}$ considered in this paper. Our description closely follows that in Pederson and Burke [28]. Each exact condition has a corresponding local condition such that if the DFA $\tilde{\epsilon}_{\text{xc}}$ satisfies the local condition, then the functional $\widetilde{E}_{\rm xc}$ satisfies the (global) exact condition.
While the exact conditions take as input the global energy $E_{\rm xc}[n(r)]$ of the DFA, the local conditions take the exchange (correlation) enhancement factor $\widetilde{F}_{\text{xc}}$, which is a function of the local value ${\epsilon}_{\text{xc}}\big{(}n(r)\big{)}$ of the exchange (correlation) energy of the DFA. For GGA functionals, the inputs $n(r)$ and $\nabla n(r)$ are usually expressed in terms of the Wigner-Seitz radius $r_{s}=(r\pi n/3)^{-1/3}$ and $s=|\nabla n|/(2(3\pi^{2})^{1/3}n^{4/3})$. Given the expressions for the exchange and correlation energies $\tilde{\epsilon}_{\text{x}}$ and $\tilde{\epsilon}_{\text{c}}$ of the DFA in terms of $s$ and $r_{s}$, we can compute $\widetilde{F}_{\text{x}}$ and $\widetilde{F}_{\text{c}}$ of $\widetilde{F}_{\text{xc}}$:
@@
  • The correlation energy $E_{\rm c}$ non-positivity condition [28] is defined as
    $E_{\rm c}[n]\leq 0.$ (3)
    The corresponding local condition is
@@
  • The $E_{\rm c}$ scaling inequality [34, 28],
    $(\gamma-1)E_{\rm c}[n_{\gamma}]\geq\gamma(\gamma-1)E_{\rm c}[n],$ (5)
    has the local condition
@@
  • The $U_{\text{c}}(\lambda)$ monotonicity condition [35, 28] is
    $\frac{dU_{\text{c}}(\lambda)}{d\lambda}\leq 0,$ (7)
    where $U_{\text{c}}(\lambda)=d(\lambda^{2}E_{\rm c}[n_{1/\lambda}])/d\lambda$ represents the correlation energy adiabatic connection curves. The corresponding local condition is
@@
  • Lieb-Oxford bound [36, 28] is
    $U_{\text{xc}}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (9)
    where $U_{\text{xc}}[n]=E_{\rm xc}[n]-T_{\text{c}}[n]$ is the potential correlation energy and $C_{\text{LO}}=2.27$ is the Lieb-Oxford constant, following [28]. The corresponding local condition is
@@
  • The Lieb-Oxford extension to $E_{\rm xc}$ [36, 28] is a generalization of the Lieb-Oxford bound with $E_{\rm xc}$ instead of $U_{XC}$:
    $E_{\rm xc}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (11)
    The corresponding local condition is
@@
  • The $T_{\text{c}}$ upper bound condition [35, 28] is
    $T_{\text{c}}[n_{\gamma}]\leq-\gamma\left(\frac{\partial E_{\rm c}[n_{\gamma}]}{\partial\gamma}\right)+E_{\rm c}[n_{\gamma}].$ (13)
    The corresponding local condition is
    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}(\infty)-\widetilde{F}_{\text{c}}}{r_{s}}.$ (14)
    where $\widetilde{F}_{\text{c}}(\infty)$ is the the limit of $\widetilde{F}_{\text{c}}$ as $r_{s}\to\infty$.
  • The conjectured $T_{\text{c}}$ upper bound [37, 38, 39, 28] is
    $T_{\text{c}}[n]\leq-E_{\rm c}[n]$ (15)
    with local condition
@@ III XCVerifier
This section describes the design the XCVerifier tool that verifies whether an DFA implemented in the Libxc library [18] satisfies the DFT exact conditions. XCVerifier consists of two main parts: XCEncoder, which encodes the given exact condition for a given Libxc functional into a formula $\psi$, and Verifier, which verifies whether this encoded formula $\psi$ is always true (valid) in the given input domain. Because DFAs involve non-linear arithmetic as well as transcendental functions, we chose the dReal solver [30] as the core solver in the Verifier. As a consequence, the formula $\psi$ generated by XCEncoder is a dReal formula.
@@ III-A XCEncoder
Given a Libxc functional, XCEncoder first translates the Maple code for $\tilde{\epsilon}_{\text{xc}}$ to Python using CodeGeneration package of Maple. To test that this translation is correct, we uniformly sample points from the input region and assert that the outputs of the translated Python implementation and the original Libxc implementation are close by a factor of at least $10^{-13}$. Of the DFAs we use in our experiments, all points for PBE, AM05, and VWN RPA are within $10^{-14}$ of the Libxc implementation. LYP is within $10^{-13}$ and SCAN is within $10^{-14}$ for $98\%$ of the points sampled.
The next step in the XCEncoder is to convert the Python code into a dReal expression. We implemented a symbolic execution engine for (a subset of) Python that generates this dReal expression. Note that though DFA implementations do not contain loops, arrays, etc., they do contain (non-recursive) function calls and if-then-else statements.
XCEncoder needs to return the dReal formula $\psi$ that encodes the given exact condition for the particular functional. Encoding the local condition corresponding to the $E_{\rm c}$ non-positivity condition is straightforward: XCEncoder uses the dReal expression for $\tilde{\epsilon}_{\text{c}}$ to directly construct the dReal formula $\tilde{\epsilon}_{\text{c}}\leq 0$.
However, the local conditions corresponding to exact conditions such as $E_{\rm c}$ scaling, Lieb-Oxford, $U_{\text{c}}$ monotonicity, and $T_{\text{c}}$ upper bound require computation of one or more derivatives, In such cases, XCEncoder uses the symbolic differentiation built into dReal to compute the derivatives. Furthermore, the local condition corresponding the $T_{\text{c}}$ upper bound condition requires computing a limit. In such cases, XCEncoder uses SymPy [31] to compute the limit, then translates this limit back to a dReal expression. After any such derivatives and limits are symbolically computed, XCEncoder constructs the required formula $\psi$ corresponding to the specific local condition.

### v5

# Towards Verifying Exact Conditions for Implementations of Density Functional Approximations

@@ Abstract
This paper presents XCVerifier, the first approach for verifying whether a DFA implementation satisfies the DFT exact conditions. XCVerifier was evaluated on five DFAs from the popular Libxc library and seven exact conditions used in recent work by Pederson and Burke. XCVerifier was able to verify or find violations for a majority of the DFA-condition pairs, demonstrating the feasibility of using formal methods to verify DFA implementations. However, it timed out on all conditions of the recent SCAN functional, revealing directions for future work.
@@ I Introduction
Density functional theory (DFT) [1, 2, 3] is a widely used approximation in the computation of electronic properties of matter. DFT calculations have been used to predict diverse properties from atomic binding energies and chemical reactivities to electronic conductivities and magnetic phenomena [4]. Consequently, DFT finds applications in a number of scientific and engineering fields, including solid state physics, computational chemistry, and materials science, and has been implemented in widely-used scientific software such as ABINIT [5], cp2k [6], ERKALE [7], Psi4 [8], Octopus [9], Qbox [10], and Quantum ESPRESSO [11].
Originating in the foundational work of Hohenberg and Kohn in 1964 [12], DFT has provided a practical way to reduce the complexity of conventional electronic structure methods by identifying the electron density function $n(r)$—a real function in three-dimensional space—as the fundamental quantity from which all other properties of a physical system can be derived. The Kohn-Sham (KS) [13] approach is currently the basis of most DFT calculations, and states that only the exchange-correlation (XC) energy portion of a functional needs to be approximated. However, the exact expression for this XC energy functional $E_{\rm xc}[n]$, which describes the complex electron-electron interactions within systems, is not known and is incredibly difficult to approximate [14].
@@
This proliferation of DFAs is reflected in the fact that the popular Libxc software library [18], which provides numerical implementations of DFAs, currently includes over 500 functionals [19]. DFAs are of varying degrees of complexity [20]: (i) local density approximations (LDAs) depend only on the electron density $n$, (ii) generalized-gradient approximations (GGAs) [21] depend on the electron density $n$ and its gradient $\nabla n$, and (iii) meta-GGAs have a further dependence on the Laplacian $\nabla^{2}n$ and the local kinetic energy density $\tau$.
@@
$\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}(\rho,\sigma)=\frac{2.884\rho^{1/3}\left(-28.944\pi^{4/3}\rho^{8/3}-0.174\pi^{2}\sigma\right)}{\pi^{1/3}\cdot\left(77.184\pi^{4/3}\rho^{8/3}+0.257\pi^{2}\sigma\right)}$
@@
However, the correlation part of PBE is significantly more complex with over 300 operations in the Libxc implementation. The SCAN meta-GGA functional [24] is even more complex with over 1000 operations, including transcendental functions such as $\exp$ and $\log$.
@@
Creating a new DFA is an art mastered by only a few researchers as of today, and the functional forms used to define the DFAs vary considerably. DFA designs fall into two categories: empirical and non-empirical. Empirically-designed DFAs (e.g., LYP [25]) are tailored for molecular chemistry applications and perform well on molecular benchmarks [26, 27]. Non-empirically designed DFAs are constructed to satisfy some exact conditions, which are known analytical properties of the exact functional (Section II). For example, the correlation energy $E_{C}$ non-positivity condition states that the correlation energy cannot be positive, i.e., $E_{\rm c}[n]\leq 0$ [28]. Furthermore, so-called norms are imposed on DFAs by requiring that they reproduce correctly some known physical systems, e.g., a hydrogen or a helium atom for which exact results are available. The SCAN functional is built to satisfy as many as 17 constraints and norms [24].
@@
Recently, Pederson and Burke (PB) [28] checked whether the Libxc implementations of various DFAs satisfy DFT exact conditions. In particular, they checked seven exact conditions by considering their corresponding local condition (Section II). These local conditions were assessed for a given DFA by performing a grid search over the inputs to the DFA and checking whether each input-output pair satisfies the local condition. Many of the conditions require gradient calculations, which were numerically approximated. The PB approach is the state of the art in the DFT community, and was the first to perform a large scale study of the role of exact conditions in density functional development. For instance, they found that many empirical DFAs satisfy these exact conditions in certain regions even though they were designed without explicit adherence to these exact conditions.
This paper addresses the problem of automatically verifying whether the implementation of a DFA satisfies the exact conditions of the density functional. It is the first to apply formal-methods techniques to density functional theory. The aim is to provide formal guarantees related to the correctness of existing DFA implementations; viz., to formally verify if the DFA implementation satisfies the exact conditions, and to determine the areas of the input domain where it does not. As a step towards solving this problem, we designed and implemented XCVerifier, a tool that verifies whether a Libxc functional, implemented in Maple[29], satisfies the given local condition (and, hence, the corresponding exact condition) using the dReal solver [30] (Section III). XCVerifier also computes any required derivatives symbolically, avoiding any potential issues arising from their numerical approximation. We also implemented a domain-splitting technique to improve the performance of XCVerifier (Algorithm 1).
@@
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke) for five popular DFAs: PBE [22], SCAN [24], LYP [25], AM05 [31], and VWN RPA [32], which cover the different types of DFAs (LDA, GGA, and meta-GGA), as well as different design categories (empirical and non-empirical). Some conditions do not apply to certain DFAs, which left us with 31 DFA-condition pairs. As shown in Section IV-B, XCVerifier was successfully able to verify or find counterexamples for 13 pairs, and it is able to partially verify an additional seven pairs (Table I). These results demonstrate the feasibility of using formal-methods techniques to verify DFA implementations. However, XCVerifier times out for 11 pairs: one property for the PBE DFA, three properties for AM05, and all of the properties for SCAN. This motivates further research on formal methods for DFT.
To further validate our approach, we compared the results from the PB approach with XCVerifier (Section IV-C): both approaches find similar regions where the conditions were violated or satisfied (Table II).
@@
Our preliminary results demonstrate the feasibility of using formal methods to prove the correctness DFT implementations, and reveal avenues for future work (Section VI).
@@ II Exact Conditions in DFT
This section lists the exact conditions of the density functional $E_{\rm xc}$ considered in this paper. Our description closely follows that in Pederson and Burke [28]. Each exact condition has a corresponding local condition such that if the DFA $\tilde{\epsilon}_{\text{xc}}$ satisfies the local condition, then the functional $\widetilde{E}_{\rm xc}$ satisfies the (global) exact condition. Note that the converse is not true: violating the local condition does not imply that the exact condition is violated. Furthermore, the region where the local condition is violated for a DFA can depend on the particular implementation of the functional. In this paper, we focus on verifying the Libxc implementation of a DFA.
The local conditions take the exchange (correlation) enhancement factor $\widetilde{F}_{\text{xc}}$, which is a function of ${\epsilon}_{\text{xc}}\big{(}n(r)\big{)}$, the local value of the exchange (correlation) energy for the DFA. Here, $n(r)$ is the electron density at the point $r$ representing the position of an electron. For GGA functionals, the inputs $n(r)$ and $\nabla n(r)$ are usually expressed in terms of the Wigner-Seitz radius $r_{s}=(4\pi n/3)^{-1/3}$ and $s=|\nabla n|/(2(3\pi^{2})^{1/3}n^{4/3})$. Given the expressions for the DFA exchange and correlation energies $\tilde{\epsilon}_{\text{x}}$ and $\tilde{\epsilon}_{\text{c}}$ in terms of $s$ and $r_{s}$, we can compute $\widetilde{F}_{\text{x}}$ and $\widetilde{F}_{\text{c}}$ of $\widetilde{F}_{\text{xc}}$:
@@
(EC1) The correlation energy ($E_{\rm c}$) non-positivity condition [28] is defined as $E_{\rm c}[n]\leq 0$. The corresponding local condition is
@@
It can also be expressed in terms of $\widetilde{F}_{\text{c}}$ as
$\widetilde{F}_{\text{c}}\geq 0.$ (4)
(EC2) The $E_{\rm c}$ scaling inequality [33, 28], $(\gamma-1)E_{\rm c}[n_{\gamma}]\geq\gamma(\gamma-1)E_{\rm c}[n]$, has the local condition
@@
(EC3) The $U_{\text{c}}(\lambda)$ monotonicity condition [34, 28] is $\frac{dU_{\text{c}}(\lambda)}{d\lambda}\leq 0$, where $U_{\text{c}}(\lambda)=d(\lambda^{2}E_{\rm c}[n_{1/\lambda}])/d\lambda$ represents the correlation energy adiabatic connection curves. The corresponding local condition is
@@
(EC4) Lieb-Oxford bound [35, 28] is $U_{\text{xc}}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}$, where $U_{\text{xc}}[n]=E_{\rm xc}[n]-T_{\text{c}}[n]$ is the potential correlation energy and $C_{\text{LO}}=2.27$ is the Lieb-Oxford constant, following [28]. The corresponding local condition is
@@
(EC5) The Lieb-Oxford extension to $E_{\rm xc}$ [35, 28] is a generalization of the Lieb-Oxford bound with $E_{\rm xc}$ instead of $U_{\text{xc}}$ :  $E_{\rm xc}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}$. The corresponding local condition is
@@
(EC6) The $T_{\text{c}}$ upper bound condition [34, 28] is $T_{\text{c}}[n_{\gamma}]\leq-\gamma\left(\frac{\partial E_{\rm c}[n_{\gamma}]}{\partial\gamma}\right)+E_{\rm c}[n_{\gamma}]$, with corresponding local condition
$\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}(\infty)-\widetilde{F}_{\text{c}}}{r_{s}}$ (9)
where $\widetilde{F}_{\text{c}}(\infty)$ is the limit of $\widetilde{F}_{\text{c}}$ as $r_{s}\to\infty$.
(EC7) The conjectured $T_{\text{c}}$ upper bound [36, 37, 38, 28] is $T_{\text{c}}[n]\leq-E_{\rm c}[n]$ with local condition
@@ III XCVerifier
This section describes the design of the XCVerifier tool, which verifies whether a DFA implemented in the Libxc library [18] satisfies the DFT local conditions in Section II. XCVerifier consists of (i) XCEncoder, which encodes the given local condition for a given Libxc functional into a formula $\psi$, and (ii) Verifier, which verifies whether this encoded formula $\psi$ is always true (valid) in the given input domain. Because DFAs involve non-linear arithmetic and transcendental functions, we chose the dReal solver [30] as the core solver in Verifier. As a consequence, the formula $\psi$ generated by XCEncoder is a dReal formula.
@@ III-A XCEncoder
Given a Libxc functional, XCEncoder first translates the Maple code for $\tilde{\epsilon}_{\text{xc}}$ to Python using the CodeGeneration package from Maple. We implemented a symbolic execution engine for (a subset of) Python that generates the dReal expression corresponding to the DFA. Though DFA implementations do not contain loops, arrays, etc., they do contain (non-recursive) function calls and if-then-else statements.
XCEncoder then constructs the dReal formula $\psi$ that encodes the given local condition for the particular functional. Encoding the local condition corresponding to the $E_{\rm c}$ non-positivity condition is straightforward: XCEncoder uses the dReal expression for $\tilde{\epsilon}_{\text{c}}$ to directly construct the dReal formula $\tilde{\epsilon}_{\text{c}}\leq 0$.
However, the local conditions corresponding to exact conditions such as $E_{\rm c}$ scaling, Lieb-Oxford, $U_{\text{c}}$ monotonicity, and $T_{\text{c}}$ upper bound require computation of one or more derivatives, In such cases, XCEncoder uses SymPy [39] to symbolically compute the derivatives. Furthermore, the local condition corresponding to the $T_{\text{c}}$ upper bound condition (Equation 9) requires computing $\lim_{r_{s}\to\infty}\widetilde{F}_{\text{c}}$. Following [28], XCEncoder substitutes an appropriately large value to approximate this limit at infinity, viz. $\widetilde{F}_{\text{c}}|_{r_{s}=100}$.